# Improving Sequential Market Clearing via Value-oriented Renewable Energy Forecasting

### v1

@@ III Methodology: Parameter Estimation
In this section, a bilevel program [18] is formulated for estimating the forecast model parameter at the training phase. Let $g(\ \cdot\ ;\Theta)$ denote the forecast model with the parameter $\Theta$, and $\bm{s}_{d,\tau}$ denote the context. The RES forecast for the time-slot $\tau$ on day $d$ is issued by,
@@
where the upper-level objective (14a) seeks to minimize the expected overall operation cost of the two markets. This is achieved by leveraging the optimal DA and RT cost functions, which are informed by the decisions obtained from the lower level (14d). (14c) limits the forecast $\hat{\bm{y}}_{d,\tau}$ within $\bar{\bm{y}}_{d,\tau}$, which can be RES capacity or its offering quantity. The lower level treats the forecast $\hat{\bm{y}}_{d,\tau}$ as an input parameter. As a consequence, both DA and RT decisions are affected by it.
To show the impact of the forecast on the operation cost more clearly, we replace the lower level with the dual problems. The overall operation cost within the upper-level objective is then substituted with the DA and RT dual objectives. These objectives are constructed as a linear combination of the right-side parameters and the associated dual variables, i.e,
@@
$\bm{\nu}_{d,1}^{*}=\mathop{\arg\max}_{\nu_{d,1}\geq 0}-\bm{\nu}_{d,1}^{\top}(\bm{\psi}_{\text{RT},d,1}+\bm{F}_{\text{RT}}\bm{x}_{d,1}^{*}),$
$\forall\tau=1,\forall d=1,...,D$ (15f)
(11)$,\forall\tau=1,\forall d=1,...,D$ (15g)
@@
At the operation phase, the forecast $\bm{\hat{y}}_{d,\tau}$ under the context $\bm{s}_{d,\tau}$ can be obtained by the trained model. Subsequently, utilizing this forecast, the operator proceeds to solve the DA market clearing in (1), followed by the RT market clearing in (2) and (3).
@@ IV Loss Function Design
We derive the loss function based on the bilevel program in (15). We first analyze how forecasts influence the DA and RT optimal solutions. Based on it, a loss function is derived, and will be used for training a forecast model.
@@ IV-A The Impact of the Forecast on the DA and RT Solutions
This section analyzes the impact of the forecasts on the primal and dual solutions of DA and RT market clearings. Specifically, we derive analytical functions that quantitatively depict the impact of the forecasts on the optimal primal and dual solutions. As a parameter to the DA market clearing (10), forecasts $\hat{\bm{y}}_{d}$ influence DA primal and dual solutions directly. As for the RT market clearing, the impact of the forecast is indirect, as it does not directly appear as the parameter in (11) or (12). Concretely, for the RT market clearing at time-slot $\tau=1$, the forecast $\hat{\bm{y}}_{d}$ influences $\bm{x}_{d,1}^{*}$, and then $\bm{x}_{d,1}^{*}$ influences the RT solutions, with the first impact being determined by the DA clearing (10), and the second impact by the RT clearing (11). For the RT clearing at time-slot $\tau=2,...,T$, the impact of the forecast is complex. The forecast $\hat{\bm{y}}_{d}$ affects the DA solutions $\bm{x}_{d,\tau}^{*},\bm{p}_{d,\tau-1}^{*}$ through the DA clearing (10). Its influence on the RT solutions $\bm{p}_{d,\tau-1}^{+-*}$ at the previous time-slot $\tau-1$ occurs as explained earlier. Then, the influence of the parameters $\bm{x}_{d,\tau}^{*},\bm{p}^{*}_{d,\tau-1},\bm{p}^{+-*}_{d,\tau-1}$ on the RT clearing solutions at time $\tau$ is determined by (12). The above process is summarized in Fig. 3.
@@
Theorem 1 implies that in a neighborhood of the parameter, the primal decision policy is represented by a linear function, whereas the corresponding dual decision policy remains a constant function. Given a specific value of $\bm{w}$, we study the local policies defined in its neighborhood. The local dual decision policy can be obtained easily, as it is a constant function. Its output is the optimal dual solution obtained by solving the dual problem of (16), given the value of $\bm{\omega}$. Additionally, after solving (16), the active constraints of (16b) can be obtained. Let $\mathcal{J}$ denote the row index set associated with (16b), and $\mathcal{J}^{a}$ denote the row index set of the active constraints. The parameters associated with the active constraints are denoted as $\bm{G}_{\mathcal{J}^{a}},\bm{\psi}_{\mathcal{J}^{a}},\bm{F}_{\mathcal{J}^{a}}$. They are the sub-matrices and sub-vectors of $\bm{G},\bm{\psi},\bm{F}$ and are comprised by the rows of $\bm{G},\bm{\psi},\bm{F}$ in the row index sets $\mathcal{J}^{a}$. We have the following proposition for the local primal decision policy,
@@
$\bm{x}^{*}=\bm{G}_{\mathcal{J}^{a}}^{-1}(\bm{\psi}_{\mathcal{J}^{a}}+\bm{F}_{\mathcal{J}^{a}}\bm{\omega})$ (17)
@@
With the active constraints of (16), we have $\bm{G}_{\mathcal{J}^{a}}\bm{x}^{*}=\bm{\psi}_{\mathcal{J}^{a}}+\bm{F}_{\mathcal{J}^{a}}\bm{\omega}$. $\bm{G}_{\mathcal{J}^{a}}^{-1}$ is the pseudo inverse of $\bm{G}_{\mathcal{J}^{a}}$. By moving $\bm{G}_{\mathcal{J}^{a}}$ from left to right, we have (17). ∎
@@
(17) calculates the inverse of $\bm{G}_{\mathcal{J}^{a}}$, whose computational complexity depends on the matrix size, i.e., $|\mathcal{J}^{a}|$. We note that a similar matrix inverse is also involved in [15], with the matrix size of $|\mathcal{J}|+\iota$, where $\iota$ is the dimension of $\bm{x}$. There is $|\mathcal{J}|+\iota>|\mathcal{J}^{a}|$. Therefore, the computation burden of [15] is larger.
With Proposition 1, we present the local primal decision policy for the DA clearing (10), and RT clearing (11) and (12). Let $\mathcal{J}^{a}_{\text{DA},d},\mathcal{J}^{a}_{\text{RT},d,\tau},\forall\tau=1,...,T$ denote the row index set of active constraints (10b), (11b), and (12b). The parameters associated with the active constraints are denoted as $\bm{G}_{\text{DA},\mathcal{J}_{\text{DA},d}^{a}},\bm{\psi}_{\text{DA},d,\mathcal{J}_{\text{DA},d}^{a}},\bm{F}_{\text{DA},\mathcal{J}_{\text{DA},d}^{a}}^{y}$, $\bm{G}_{\text{RT},\mathcal{J}_{\text{RT},d,\tau}^{a}},\bm{\psi}_{\text{RT},d,\tau,\mathcal{J}_{\text{RT},d,\tau}^{a}},\bm{F}_{\text{RT},\mathcal{J}_{\text{RT},d,\tau}^{a}}$, and $\bm{G}_{\text{RT},\mathcal{J}_{\text{RT},d,\tau}^{a}}^{\prime},\bm{\psi}_{\text{RT},d,\tau,\mathcal{J}_{\text{RT},d,\tau}^{a}}^{\prime},\bm{F}_{\text{RT},\mathcal{J}_{\text{RT},d,\tau}^{a}}^{\prime x},\bm{F}_{\text{RT},\mathcal{J}_{\text{RT},d,\tau}^{a}}^{\prime p},\bm{F}_{\text{RT},\mathcal{J}_{\text{RT},d,\tau}^{a}}^{\prime+-}$, respectively. We have the following proposition for the local primal decision policies of (10), (11), and (12).
@@
$\bm{x}_{d}^{*}=\bm{G}_{\text{DA},\mathcal{J}_{\text{DA},d}^{a}}^{-1}(\bm{\psi}_{\text{DA},d,\mathcal{J}_{\text{DA},d}^{a}}+\bm{F}_{\text{DA},\mathcal{J}_{\text{DA},d}^{a}}^{y}\hat{\bm{y}}_{d})$ (18)
$\bm{z}_{d,\tau}^{*}=\bm{G}_{\text{RT},\mathcal{J}_{\text{RT},d,\tau}^{a}}^{-1}(\bm{\psi}_{\text{RT},d,\tau,\mathcal{J}_{\text{RT},d,\tau}^{a}}+\bm{F}_{\text{RT},\mathcal{J}^{a}_{\text{RT},d,\tau}}\bm{x}_{d,\tau}^{*}),\forall\tau=1$ (19)
$\begin{split}&\bm{z}_{d,\tau}^{*}=\bm{G}_{\text{RT},\mathcal{J}_{\text{RT},d,\tau}^{a}}^{\prime-1}(\bm{\psi}^{\prime}_{\text{RT},d,\tau,\mathcal{J}_{\text{RT},d,\tau}^{a}}+\bm{F}_{\text{RT},\mathcal{J}_{\text{RT},d,\tau}^{a}}^{\prime x}\bm{x}_{d,\tau}^{*}+\\ &\bm{F}_{\text{RT},\mathcal{J}_{\text{RT},d,\tau}^{a}}^{\prime p}\bm{p}_{d,\tau-1}^{*}+\bm{F}_{\text{RT},\mathcal{J}_{\text{RT},d,\tau}^{a}}^{\prime+-}\bm{p}_{d,\tau-1}^{+-*}),\forall\tau=2,...,T\end{split}$ (20)
Eq. (18) is a linear function of $\hat{\bm{y}}_{d}$, whose output is the DA solution $\bm{x}_{d}^{*}$ on the day $d$. In the following, we show how to rewrite (19) and (20) as a function of the forecast $\hat{\bm{y}}_{d}$.
In addition, we define an operator $\Pi_{\mathcal{J}}:h(\bm{x})\mapsto\tilde{h}(\bm{x})$, where $h(\bm{x})=\bm{A}\bm{x}+\bm{b}$ is a linear function with the parameter $\bm{A},\bm{b}$, and $\mathcal{J}$ is the row index subset of $\bm{A},\bm{b}$. The output $\tilde{h}(\bm{x})$ of the operator is also a linear function, where $\tilde{h}(\bm{x})=\bm{A}_{\mathcal{J}}\bm{x}+\bm{b}_{\mathcal{J}}$ and $\bm{A}_{\mathcal{J}},\bm{b}_{\mathcal{J}}$ are the sub-matrix and sub-vector of $\bm{A},\bm{b}$.
@@
$\begin{split}&f^{x}_{d,\tau}(\hat{\bm{y}}_{d}):=\bm{x}_{d,\tau}^{*}=\Pi_{\mathcal{I}_{\text{DA},d,\tau}}(\bm{G}_{\text{DA},\mathcal{J}_{\text{DA},d}^{a}}^{-1}\bm{F}_{\text{DA},\mathcal{J}_{\text{DA},d}^{a}}^{y}\hat{\bm{y}}_{d}+\\ &\bm{G}_{\text{DA},\mathcal{J}_{\text{DA},d}^{a}}^{-1}\bm{\psi}_{\text{DA},d,\mathcal{J}_{\text{DA},d}^{a}}),\forall\tau=1,...,T\end{split}$ (21)
The coefficients are determined by the coefficients of (18), whose row indexes belong to the set $\mathcal{I}_{\text{DA},d,\tau}$. By substituting (21) into (19), we can obtain the function between RT primal solution $\bm{z}_{d,\tau}^{*}$ at time-slot $\tau=1$ and the forecast $\hat{\bm{y}}_{d}$.
$\begin{split}&f_{d,\tau}^{z}(\hat{\bm{y}}_{d}):=\bm{z}_{d,\tau}^{*}=\bm{G}_{\text{RT},\mathcal{J}_{\text{RT},d,\tau}^{a}}^{-1}\bm{F}_{\text{RT},\mathcal{J}^{a}_{\text{RT},d,\tau}}f_{d,\tau}^{x}(\hat{\bm{y}}_{d})+\\ &\bm{G}_{\text{RT},\mathcal{J}_{\text{RT},d,\tau}^{a}}^{-1}\bm{\psi}_{\text{RT},d,\tau,\mathcal{J}_{\text{RT},d,\tau}^{a}},\forall\tau=1\end{split}$ (22)
@@
Likewise, since $\bm{p}_{d,\tau-1}^{+-*}$ is a part of $\bm{z}_{d,\tau-1}^{*}$, let $\mathcal{I}_{\text{RT},d,\tau-1}^{+-}$ be the row index set corresponds to $\bm{p}_{d,\tau-1}^{+-*}$ within $\bm{z}_{d,\tau-1}^{*}$. We can express a function of $\bm{p}_{d,\tau-1}^{+-*}$ w.r.t. $\hat{\bm{y}}_{d}$ as,
@@
$\begin{split}&f^{z}_{d,\tau}(\hat{\bm{y}}_{d}):=\\ &\bm{z}_{d,\tau}^{*}=\bm{G}_{\text{RT},\mathcal{J}_{\text{RT},d,\tau}^{a}}^{\prime-1}(\bm{\psi}^{\prime}_{\text{RT},d,\tau,\mathcal{J}_{\text{RT},d,\tau}^{a}}+\bm{F}_{\text{RT},\mathcal{J}_{\text{RT},d,\tau}^{a}}^{\prime x}f_{d,\tau}^{x}(\hat{\bm{y}}_{d})+\\ &\bm{F}_{\text{RT},\mathcal{J}_{\text{RT},d,\tau}^{a}}^{\prime p}f_{d,\tau-1}^{p}(\hat{\bm{y}}_{d})+\bm{F}_{\text{RT},\mathcal{J}_{\text{RT},d,\tau}^{a}}^{\prime+-}f_{d,\tau-1}^{+-}(\hat{\bm{y}}_{d})),\\ &\forall\tau=2,...,T\end{split}$ (25)
To sum up, the function between the forecast $\hat{\bm{y}}_{d}$ and DA and RT primal solutions, which is defined in the neighborhood of $\hat{\bm{y}}_{d}$, are summarized in the (21)-(25). The function between the optimal dual solutions and the forecast $\hat{\bm{y}}_{d}$ is a constant function, whose output can be obtained by solving the dual problems of (10),(11), and (12). With these, we are ready to transform the upper-level objective (15a) to a function of the forecast $\hat{\bm{y}}_{d}$. The details are in the next subsection.

### v2

@@ III Methodology: Parameter Estimation
In this section, we show how to identify the RIEQ, whose value corresponds to $\hat{\bm{y}}_{d,\tau}$, an input parameter in (1), via training forecasting models. At the training phase, a bilevel program [18] is formulated for estimating the forecasting model parameter. Let $g(\ \cdot\ ;\Theta)$ denote the forecasting model with the parameter $\Theta$, and $\bm{s}_{d,\tau}$ denote the context. The RIEQ forecast for the time $\tau$ on day $d$ is,
@@
where the upper-level objective (14a) seeks to minimize the expected overall operating cost of the two markets. This is achieved by leveraging the optimal DA and RT cost functions, which are informed by the decisions obtained from the lower level (14d). (14c) limits the forecast $\hat{\bm{y}}_{d,\tau}$ within $\bar{\bm{y}}_{d,\tau}$, which can be RES capacity. The lower level treats the forecast $\hat{\bm{y}}_{d,\tau}$ as an input parameter. As a consequence, both DA and RT decisions are affected by it.
To show the impact of the forecast on the operating cost more clearly, we replace the lower level with the dual problems. The overall operating cost within the upper-level objective is then substituted with the DA and RT dual objectives. These objectives are constructed as a linear combination of the right-side parameters and the associated dual variables, i.e,
@@
$\bm{\nu}_{d,1}^{*}=\mathop{\arg\max}_{\nu_{d,1}\geq 0}-\bm{\nu}_{d,1}^{\top}(\bm{\psi}_{\text{RT},d,1}+\bm{F}_{\text{RT}}\bm{x}_{d,1}^{*}),\forall d=1,...,D$ (15f)
(11)$,\forall d=1,...,D$ (15g)
@@
At the operation phase, the forecast $\bm{\hat{y}}_{d,\tau}$ under the context $\bm{s}_{d,\tau}$ can be obtained by the trained model. Subsequently, utilizing this forecast, the operator proceeds to solve the DA clearing in (1), followed by the RT clearing in (2) and (3).
@@ IV Loss Function Design
We derive the loss function based on the bilevel program in (15). We first analyze how forecasts influence the DA and RT optimal solutions. Based on it, a loss function is derived, and will be used for training a forecasting model.
@@ IV-A The Impact of the Forecast on the DA and RT Solutions
This section analyzes the impact of the forecasts on the primal and dual solutions of DA and RT market clearings. Specifically, we derive analytical functions that quantitatively depict the impact of the forecasts on the optimal primal and dual solutions. As a parameter to the DA market clearing (10), forecasts $\hat{\bm{y}}_{d}$ influence DA primal and dual solutions directly. As for the RT market clearing, the impact of the forecast is indirect, as it does not directly appear as the parameter in (11) or (12). Concretely, for the RT market clearing at time-slot $\tau=1$, the forecast $\hat{\bm{y}}_{d}$ influences $\bm{x}_{d,1}^{*}$, and then $\bm{x}_{d,1}^{*}$ influences the RT solutions, with the first impact being determined by the DA clearing (10), and the second impact by the RT clearing (11). For the RT clearing at time-slot $\tau=2,...,T$, the impact of the forecast is more complex. The forecast $\hat{\bm{y}}_{d}$ affects the DA solutions $\bm{x}_{d,\tau}^{*},\bm{p}_{d,\tau-1}^{*}$ through the DA clearing (10). The influence of the forecast $\hat{\bm{y}}_{d}$ on the RT solutions $\bm{p}_{d,\tau-1}^{+-*}$ at the previous time-slot $\tau-1$ occurs as explained earlier. Then, the influence of the parameters $\bm{x}_{d,\tau}^{*},\bm{p}^{*}_{d,\tau-1},\bm{p}^{+-*}_{d,\tau-1}$ on the RT clearing solutions at time $\tau$ is determined by (12). The above process is summarized in Fig. 3.
@@
Theorem 1 implies that in a neighborhood of the parameter, the primal decision policy is represented by a linear function, whereas the corresponding dual decision policy remains a constant function. Given a specific value of $\bm{w}$, we study the local policies defined in its neighborhood. The local dual decision policy can be obtained easily, as it is a constant function. Its output is the optimal dual solution obtained by solving the dual problem of (16), given the value of $\bm{\omega}$. Additionally, after solving (16), the active constraints of (16b) can be obtained. Let $\mathcal{J}$ denote the row index set associated with (16b), and $\mathcal{J}^{a}$ denote the row index set of the active constraints. The parameters associated with the active constraints are denoted as $\bm{G}[\mathcal{J}^{a}],\bm{\psi}[\mathcal{J}^{a}],\bm{F}[\mathcal{J}^{a}]$. They are the sub-matrices and sub-vectors of $\bm{G},\bm{\psi},\bm{F}$ and are comprised by the rows of $\bm{G},\bm{\psi},\bm{F}$ in the row index sets $\mathcal{J}^{a}$. We have the following proposition for the local primal decision policy,
@@
$\bm{x}^{*}=\bm{G}[\mathcal{J}^{a}]^{-1}(\bm{\psi}[\mathcal{J}^{a}]+\bm{F}[\mathcal{J}^{a}]\bm{\omega})$ (17)
@@
With the active constraints of (16), we have $\bm{G}[\mathcal{J}^{a}]\bm{x}^{*}=\bm{\psi}[\mathcal{J}^{a}]+\bm{F}[\mathcal{J}^{a}]\bm{\omega}$. $\bm{G}[\mathcal{J}^{a}]^{-1}$ is the pseudo inverse of $\bm{G}[\mathcal{J}^{a}]$. By moving $\bm{G}[\mathcal{J}^{a}]$ from left to right, we have (17). ∎
@@
(17) calculates the inverse of $\bm{G}[\mathcal{J}^{a}]$, whose computational complexity depends on the matrix size, i.e., $|\mathcal{J}^{a}|$. We note that a similar matrix inverse is also involved in [14], with the matrix size of $|\mathcal{J}|+\iota$, where $\iota$ is the dimension of $\bm{x}$. There is $|\mathcal{J}|+\iota>|\mathcal{J}^{a}|$. Therefore, the computation burden of [14] is larger.
With Proposition 1, we present the local primal decision policy for the DA clearing (10), and RT clearing (11) and (12). Let $\mathcal{J}^{a}_{\text{DA},d},\mathcal{J}^{a}_{\text{RT},d,\tau},\forall\tau=1,...,T$ denote the row index set of active constraints (10b), (11b), and (12b). The parameters associated with the active constraints are denoted as $\bm{G}_{\text{DA}}[\mathcal{J}_{\text{DA},d}^{a}],\bm{\psi}_{\text{DA},d}[\mathcal{J}_{\text{DA},d}^{a}],\bm{F}_{\text{DA}}^{y}[\mathcal{J}_{\text{DA},d}^{a}]$, $\bm{G}_{\text{RT}}[\mathcal{J}_{\text{RT},d,\tau}^{a}],\bm{\psi}_{\text{RT},d,\tau}[\mathcal{J}_{\text{RT},d,\tau}^{a}],\bm{F}_{\text{RT}}[\mathcal{J}_{\text{RT},d,\tau}^{a}]$,$\bm{G}_{\text{RT}}^{\prime}[\mathcal{J}_{\text{RT},d,\tau}^{a}]$, and $\bm{\psi}_{\text{RT},d,\tau}^{\prime}[\mathcal{J}_{\text{RT},d,\tau}^{a}],\bm{F}_{\text{RT}}^{\prime x}[\mathcal{J}_{\text{RT},d,\tau}^{a}],\bm{F}_{\text{RT}}^{\prime p}[\mathcal{J}_{\text{RT},d,\tau}^{a}],\bm{F}_{\text{RT}}^{\prime+-}[\mathcal{J}_{\text{RT},d,\tau}^{a}]$. We have the following proposition for the local primal decision policies of (10), (11), and (12).
@@
$f^{x}_{d}(\hat{\bm{y}}_{d}):=\bm{x}_{d}^{*}=\bm{G}_{\text{DA}}[\mathcal{J}_{\text{DA},d}^{a}]^{-1}(\bm{\psi}_{\text{DA},d}[\mathcal{J}_{\text{DA},d}^{a}]+\bm{F}_{\text{DA}}^{y}[\mathcal{J}_{\text{DA},d}^{a}]\hat{\bm{y}}_{d})$ (18)
$\begin{split}&\bm{z}_{d,\tau}^{*}=\bm{G}_{\text{RT}}[\mathcal{J}_{\text{RT},d,\tau}^{a}]^{-1}(\bm{\psi}_{\text{RT},d,\tau}[\mathcal{J}_{\text{RT},d,\tau}^{a}]+\\ &\bm{F}_{\text{RT}}[\mathcal{J}^{a}_{\text{RT},d,\tau}]\bm{x}_{d,\tau}^{*}),\tau=1\end{split}$ (19)
$\begin{split}&\bm{z}_{d,\tau}^{*}=\bm{G}_{\text{RT}}^{\prime}[\mathcal{J}_{\text{RT},d,\tau}^{a}]^{-1}(\bm{\psi}^{\prime}_{\text{RT},d,\tau}[\mathcal{J}_{\text{RT},d,\tau}^{a}]+\bm{F}_{\text{RT}}^{\prime x}[\mathcal{J}_{\text{RT},d,\tau}^{a}]\bm{x}_{d,\tau}^{*}+\\ &\bm{F}_{\text{RT}}^{\prime p}[\mathcal{J}_{\text{RT},d,\tau}^{a}]\bm{p}_{d,\tau-1}^{*}+\bm{F}_{\text{RT}}^{\prime+-}[\mathcal{J}_{\text{RT},d,\tau}^{a}]\bm{p}_{d,\tau-1}^{+-*}),\forall\tau=2,...,T\end{split}$ (20)
Eq. (18) is a linear function of $\hat{\bm{y}}_{d}$, whose output is the DA solution $\bm{x}_{d}^{*}$ on the day $d$. In the following, we show how to rewrite (19) and (20) as a function of the forecast $\hat{\bm{y}}_{d}$. The illustration is in Fig. 4.
Figure 4: An illustration of the calculation for rewriting (19) for $\tau=1$ and (20) for $\tau=2$ as functions of the forecast $\hat{\bm{y}}_{d}$. Different functions are represented by boxes in various colors. The diamond represents the calculation of the operator $\Pi_{\mathcal{J}}$ to obtain the function relating part of the vector to the forecast $\bm{\hat{y}}_{d}$. The circle represents the linear transformation.
In addition, we define an operator $\Pi_{\mathcal{J}}:h(\bm{x})\mapsto\tilde{h}(\bm{x})$, where $h(\bm{x})=\bm{A}\bm{x}+\bm{b}$ is a linear function with the parameter $\bm{A},\bm{b}$, and $\mathcal{J}$ is the row index subset of $\bm{A},\bm{b}$. The output $\tilde{h}(\bm{x})$ of the operator is also a linear function, where $\tilde{h}(\bm{x})=\bm{A}[\mathcal{J}]\bm{x}+\bm{b}[\mathcal{J}]$ and $\bm{A}[\mathcal{J}],\bm{b}[\mathcal{J}]$ are the sub-matrix and sub-vector of $\bm{A},\bm{b}$.
@@
$\begin{split}&f^{x}_{d,\tau}(\hat{\bm{y}}_{d}):=\bm{x}_{d,\tau}^{*}=\Pi_{\mathcal{I}_{\text{DA},d,\tau}}(f^{x}_{d}(\hat{\bm{y}}_{d})),\forall\tau=1,...,T\end{split}$ (21)
The coefficients are determined by the coefficients of $f^{x}_{d}(\hat{\bm{y}}_{d})$ in (18), whose row indexes belong to the set $\mathcal{I}_{\text{DA},d,\tau}$. By substituting (21) into (19), we can obtain the function between RT primal solution $\bm{z}_{d,\tau}^{*}$ at time $\tau=1$ and the forecast $\hat{\bm{y}}_{d}$.
$\begin{split}&f_{d,\tau}^{z}(\hat{\bm{y}}_{d}):=\bm{z}_{d,\tau}^{*}=\bm{G}_{\text{RT}}[\mathcal{J}_{\text{RT},d,\tau}^{a}]^{-1}\bm{F}_{\text{RT}}[\mathcal{J}^{a}_{\text{RT},d,\tau}]f_{d,\tau}^{x}(\hat{\bm{y}}_{d})+\\ &\bm{G}_{\text{RT}}[\mathcal{J}_{\text{RT},d,\tau}^{a}]^{-1}\bm{\psi}_{\text{RT},d,\tau}[\mathcal{J}_{\text{RT},d,\tau}^{a}],\tau=1\end{split}$ (22)
@@
Likewise, since $\bm{p}_{d,\tau-1}^{+-*}$ is a part of $\bm{z}_{d,\tau-1}^{*}$, let $\mathcal{I}_{\text{RT},d,\tau-1}^{+-}$ be the row index set corresponds to $\bm{p}_{d,\tau-1}^{+-*}$ within $\bm{z}_{d,\tau-1}^{*}$. We can express the function of $\bm{p}_{d,\tau-1}^{+-*}$ w.r.t. $\hat{\bm{y}}_{d}$ as,
@@
$\begin{split}&f^{z}_{d,\tau}(\hat{\bm{y}}_{d}):=\\ &\bm{z}_{d,\tau}^{*}=\bm{G}_{\text{RT}}^{\prime}[\mathcal{J}_{\text{RT},d,\tau}^{a}]^{-1}(\bm{\psi}^{\prime}_{\text{RT},d,\tau}[\mathcal{J}_{\text{RT},d,\tau}^{a}]+\\ &\bm{F}_{\text{RT}}^{\prime x}[\mathcal{J}_{\text{RT},d,\tau}^{a}]f_{d,\tau}^{x}(\hat{\bm{y}}_{d})+\bm{F}_{\text{RT}}^{\prime p}[\mathcal{J}_{\text{RT},d,\tau}^{a}]f_{d,\tau-1}^{p}(\hat{\bm{y}}_{d})+\\ &\bm{F}_{\text{RT}}^{\prime+-}[\mathcal{J}_{\text{RT},d,\tau}^{a}]f_{d,\tau-1}^{+-}(\hat{\bm{y}}_{d})),\forall\tau=2,...,T\end{split}$ (25)
To sum up, the linear functions between the forecast $\hat{\bm{y}}_{d}$ and DA and RT primal solutions, which are defined in the neighborhood of $\hat{\bm{y}}_{d}$, are summarized in the (21)-(25). The function between the optimal dual solutions and the forecast $\hat{\bm{y}}_{d}$ is a constant function, whose output can be obtained by solving the dual problems of (10),(11), and (12). With these, we are ready to transform the upper-level objective (15a) to a function of the forecast $\hat{\bm{y}}_{d}$. The details are in the next subsection.